OpenStudy (anonymous):

Let P(x) be a polynomial of degree 2m. Show that \( P(x^2)= P(x)P(x-1) \) if and only if \[ P(x) =\left( x^2 + x+1 \right)^m\]

OpenStudy (anonymous):

See http://openstudy.com/study#/updates/4ff6b092e4b01c7be8c98dee for the particular case where m=5.

OpenStudy (anonymous):

hmm looks like all the heavy lifting was done showing that since the only roots are \(-\frac{1}{2}\pm\frac{\sqrt{3}}{2}i\) it must be \((x^2+x+1)^m\) is that the hook, or is there something else?

OpenStudy (anonymous):

I agree. I posted so people can see that it is true in this general form.

OpenStudy (anonymous):

ok then kudos to @mukushla who did all the work for this one